# Restless Bandit Problem with Rewards Generated by a Linear Gaussian Dynamical System

### v1

@@ C.3 Regret Performance
Regret satisfies the following inequality with a probability of at least $\left(1-\delta\right)^{4}$
@@
$\phi_{a,a^{*}}^{2}$ $\triangleq u_{a^{*}}-u_{a}+\sigma\left(\delta,c_{a^{*}}\mid\mathbf{c}\right)+\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)$ (45)
@@
and $\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)$ is defined to be (32).
@@
$G_{c_{a^{*}}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)+u_{a^{*}}+\sigma\left(\delta,c_{a^{*}}\mid\mathbf{c}\right)\geq G_{c_{a}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)+u_{a}-\sigma\left(\delta,c_{a}\mid\mathbf{c}\right).$
@@
$\left(G_{c_{a^{*}}\mid\mathbf{c}}-G_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi\left(\mathbf{c}\right)\geq u_{a}-\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)-u_{a^{*}}-\sigma\left(\delta,c_{a^{*}}\mid\mathbf{c}\right).$
@@
The regret bound is a function of prediction errors $\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)$ and $\sigma\left(\delta,c_{a^{*}}\mid\mathbf{c}\right)$. To control the impact of prediction error $\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)$, we set $u_{a}$ and $u_{a^{*}}$ to be $u_{a}=\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)$ and $u_{a^{*}}=\sigma\left(\delta,c_{a^{*}}\mid\mathbf{c}\right)$. Therefore, by setting $u_{a}=\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)$ for all $c_{a}\in\mathcal{A}$ and using Theorem 5, Theorem 2 can be easily proven.

### v2

@@ C.3 Regret Performance
Let $\delta_{e}=\delta_{b}=\delta\in(0,1)$. Regret satisfies the following inequality with a probability of at least $\left(1-\delta\right)^{4}$
@@
$\phi_{a,a^{*}}^{2}$ $\triangleq u_{a^{*}}-u_{a}+\sigma\left(\delta,\delta,c_{a^{*}}\mid\mathbf{c}\right)+\sigma\left(\delta,\delta,c_{a}\mid\mathbf{c}\right)$ (44)
@@
and $\sigma\left(\delta,\delta,c_{a}\mid\mathbf{c}\right)$ is defined to be (32).
@@
$G_{c_{a^{*}}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)+u_{a^{*}}+\sigma\left(\delta,\delta,c_{a^{*}}\mid\mathbf{c}\right)\geq G_{c_{a}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)+u_{a}-\sigma\left(\delta,\delta,c_{a}\mid\mathbf{c}\right).$
@@
$\left(G_{c_{a^{*}}\mid\mathbf{c}}-G_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi\left(\mathbf{c}\right)\geq u_{a}-\sigma\left(\delta,\delta,c_{a}\mid\mathbf{c}\right)-u_{a^{*}}-\sigma\left(\delta,\delta,c_{a^{*}}\mid\mathbf{c}\right).$
@@
The regret bound is a function of prediction errors $\sigma\left(\delta,\delta,c_{a}\mid\mathbf{c}\right)$ and $\sigma\left(\delta,\delta,c_{a^{*}}\mid\mathbf{c}\right)$. To control the impact of prediction error $\sigma\left(\delta,\delta,c_{a}\mid\mathbf{c}\right)$, we set $u_{a}$ and $u_{a^{*}}$ to be $u_{a}=\sigma\left(\delta,\delta,c_{a}\mid\mathbf{c}\right)$ and $u_{a^{*}}=\sigma\left(\delta,\delta,c_{a^{*}}\mid\mathbf{c}\right)$. By setting $u_{a}=\sigma\left(\delta,\delta,c_{a}\mid\mathbf{c}\right)$ for all $c_{a}\in\mathcal{A}$, setting $\delta_{e}=\delta_{b}=\delta\in(0,1)$, and using Lemma 4 and Theorem 5, Theorem 2 can be easily proven.